What are the solutions of |2x+1|-3=6
Answer by stanbon(75887)

What are the solutions of |2x+1|-3=6
-----
|2x+1| = 9
2x+1 = 9 or 2x+1 = -9
2x = 8 or 2x = -10
x = 4 or x = -5
------------
Cheers,
Stan H.
==============